OpenStudy (anonymous):

Simplify e^-4 a)-4e b)-4 c)-1/4 d)1/-4e

OpenStudy (michele_laino):

applying the properties of powers, I can write: \[\large {e^{ - 4}} = {\left( {\frac{1}{e}} \right)^4} = \frac{1}{{{e^4}}}\]

OpenStudy (anonymous):

oh i wrote the problem incorrectly

OpenStudy (anonymous):

im going to redo this question

OpenStudy (michele_laino):

ok!